Multiple Choice Question 88 The per-unit standards for direct labor are 2 direct labor hours at $15 per hour. If in producing 2100 units, the actual direct labor cost was $62400 for 3900 direct labor hours worked, the total direct labor variance is $600 unfavorable. $375 unfavorable. $2100 unfavorable. O $600 favorable.
Answer #1

88) Total labour variance = (2100*2*15)-62400 = 600 Favorable

So answer is d) $600 Favorable

90) Labour price variance = (Standard rate-actual rate)actual hours

                    -10300           = (13-X)*10300

actual rate = 14 per hour

So answer is b) 14.00 per labour hour

100) Difference between standard and actual price = 2000/20000 = 0.10

So answer is b) $0.10

105) Total material variance = (5200*2)-(5400*2.2) = 1480 U

So answer is a) $1480 U